A ferris wheel has a radius of 7.62m and spins at a constant speed of 0.67m/s. If you were riding in 

one of the baskets of the ferris wheel:



What is your centripetal acceleration as you ride the ferris wheel?




1
Expert's answer
2021-10-26T09:39:55-0400
"a_c=\dfrac{v^2}{r}=\dfrac{(0.67\ \dfrac{m}{s})^2}{7.62\ m}=0.06\ \dfrac{m}{s^2}."